OpenStudy (anonymous):

If f(x) = ax^2 - 12x + c and f(-3/2) = 8 is the maximum value, find the values of a and c.

OpenStudy (***[isuru]***):

it has to do something with mentioning my name... since the username contain *** signs the question goes numb when someone mentions my name :/ anyway where was u ? did u make it a complete square ?

Parth (parthkohli):

Erm. f(x) = ax^2 - 12x + c f(-3/2) = 8 but f(-3/2) also equals a(-3/2)^2 - 12(-3/2) + c. Also note that the x-coordinate at the max value is -b/2a. Good luck!

OpenStudy (***[isuru]***):

y = ax^2 -12x + c = a [ x^2 - (12/a)x + c/a ] = a [ ( x -6/a )^2 + c/a - 36/a^2 ] = a [ ( x-6/a)^2 + (ac -36)/a^2 ] = a ( x - 6/a)^2 + (ac - 36)a did u make it this far... ?

OpenStudy (anonymous):

I thought that when you're completing a square, you only divide the first and second term by a?

Parth (parthkohli):

-b/2a is -3/2. b is -12.

OpenStudy (***[isuru]***):

yes ... but u must make it a perfect square

OpenStudy (***[isuru]***):

when this graph takes it maximum value a( x - 6)^2 should equal to zero then the max value 8 will equal to (ac - 36)/a which means (ac -36)/a = 8 and now u have another equation ...

OpenStudy (***[isuru]***):

now all u have to do is solve the 2 simultaneous equations and get the values for a , c

OpenStudy (***[isuru]***):

the first one was 9a + 4c = -40 ------(1) ans the second one is 8a - ac = -36 -----(2)

OpenStudy (***[isuru]***):

 got it ?